FEL Upgrade Project Weekly Brief –June 24-28, 2002



This was a very productive week: (1) the RF system for the second cryomodule (Zone 4) was

completely checked out except for the final SRF test when the 2nd cryomodule is installed. All 8 klystrons are operational; (2) the new High Voltage tank for the Upgrade gun arrived on site and was installed in the FEL vault; (3) all the quadrupole girder assemblies between the first arc and the optical chicane are installed;(4) we received favorable bids on our last major procurement (the large arc vacuum chambers) and expect to place the order next week.

Note: We will not be sending a weekly report out next Friday due to the 4th of July holiday.

Many of the FEL team members will be taking advantage of a long weekend as we prepare for the big push to install the remainder of the machine over the summer months.

We wish a pleasant holiday to all of our friends and collaborators.


We worked with the DOE, ONR and AFRL program offices this week to tie up the loose ends we have with remaining FY02 funding which has not yet been transferred to the DOE/SURA contract for FEL activities. We thank Gil Graff, John Eric and Wayne Skinner for their assistance in these matters.

On Thursday June 27, F.Dylla attended the quarterly meeting of the Virginia Research and Technology Advisory Commission which is chaired by RADM(ret) Marc Pelaez and Prof. Anita Jones (UVA). The VRTAC has always been very supportive of the FEL project.

We met this week with Brian Hankla of the JTO Lethality TAWG to discuss continuation of our FEL laser effects testing program on materials. Our proposal to construct a dedicated suite of materials diagnostics and to continue with beam tests has a high priority with the Navy. We expect about $600k to be provided for this effort next year.

WBS 3 (Beam Physics):

Initial (organizational) planning and detailing of the commissioning plan started. Design of a very compact, very large acceptance energy recovery system continued.





WBS 4 (Injector):

The gun stand and chamber were installed in the injector pit and is being aligned. W&M's implant chamber was cleaned to remove flaking SiO2 and the high voltage connection electrodes were polished. The ball cathode was installed into the chamber with the first implant run taking place today. Detail work on the two shielded bellows and SF6 tank continue.

The final drawings for the light box support structure were signed off this week. The drawing for the light box transport was ready for final check. The machine shop is making good progress on the light box can. The only design work remaining on the light box is the alignment drawing.

Gun HVPS - The tank for the Gun's HVPS was delivered this week and all but the door was installed. The Gun Tank must be installed before the final mounting alignment can be completed and tank's door is installed. Due to dust and dirt in the vault getting on the HV components, the HVPS will be installed after the door is in place and can be closed. The draft SOP was sent out for comments and any corrections.

WBS 6 (RF):

Zone 4 - The zone was restored to operating condition. RF was run through all 8 klystrons. The RF system for unit 4-8 was calibrated and data curves recorded. All systems look good and are ready for the cryomodule.The MPS & PSS equipment needs to be installed for this zone. Prior to SRF checkout of the cryomodule that is being installed in July, the MPS and PSS systems need to be activated and certified again.

Quarter HVPS - The PLC code to control the klystron heater and the SCR controller was successfully tested on unit 3. The printed circuit board to monitor the heater voltage was released this week for fabrication. The draft SOP is in progress and will be sent out for

comments and corrections next week.

Quarter Klystrons - Wiring of the klystron carts continue.

WBS 8 (Instrumentation):

Nate (our DOE student) has completed the gain and characterizing of the BLM's (good job!). All assemblies have been tested, packed and stored. Additionally he investigated response to various other wavelength LEDs.

Work continues on the fast raster for the recirculating beam dump. The system consists of a control chassis upstairs in the control room and a tunnel driver chassis near the magnet. Three (horizontal & vertical + spare) of the tunnel boards have been stuffed and are currently being tested. The upstairs Fast Raster board check prints have been turned in to EECAD for final sign off. The new PCBs should be ready for testing in a couple weeks.

The Canbus beam viewer interface boards have been received from the board house and the CAD drawings turned over to the EECAD. These will stuffed (surface mount) and tested then sent to Al Grippo for software debugging.

The students are working on all of the girder connections including air and water. The beam viewer assemblies, cameras and lamp rings on the girders have been installed and tested up to the viewer chassis along with the initial alignment of the cameras and mirrors in the 3F region. The rework of the video mux and distribution system has also begun this week. A considerable amount of effort has gone into planning and documenting the AMS/Video system. One goal is to have all of the 3F region viewers completely operational with video to the control room by July 4 break.

An updated version of the check print for the BLM front card has been corrected and returned to the EECAD group for final changes. This implies that the next check print should prove signable (F0151 schematic). In parallel, the BLM front cardboard layout continues to make progress. The design for the synchronous countdown carrier card is nearing completion for the upgrade to the pulse control system for the drive laser.

Machine Protection System planning is now in full swing. The design review will be held the week of July 8. A very simple and sound approach is being taken; in principle each system interfaces to a single opto-isolated VME board which intern generates allowable MACHINE MODES. These boards are then connected to a master which generates the allowable BEAM MODES. This in turn feeds the Drive Laser Pulse Controller (DLPC) which controls the electron source.

WBS 9 (Transport):


Optical Chicane Dipoles (GW)

• At Magnet Enterprises International (MEI) in Oakland CA. the first GW dipole is almost done.

• The first unit passed pre assembly with a few fastener lengths requiring change.

• MEI glued the first sets of Purcell Gap shims to two dipoles. The first was uniform through most of the area but for one .003 bump. They stripped the shims off and are trying again, this time with booties and hair nets in the enclosure during gluing. The other core’s glue-down quality is being evaluated.

• Cores for the Second batch of 6 UV Magnets remain in final machining.

• MEI used a different batch of the same epoxy for Coil #13. This did not fix the welt problem. The epoxy from the formulator that is with out flexiblizer finally reached MEI and they will pot # 14 with that material today.

Injector Dipoles (DU/DV)

• At Master Machine, The first GV core was shipped out to Wang NMR for delivery on Tuesday and the remaining 5 units are ready for gap inspection and shipment Monday. GU cores will follow.

• All GV and GU coils are potted and ready for insertion in cores at Wang NMR.

Arc 180 Degree Dipoles (GY)

• The 2nd core was received by Wang from Bosma.

• At Wang NMR, the third GY Coil (of four) is half wound.

• GY Coil Potting fixture remains in fabrication.

• Engineering Change Orders that correct some problems with the water fittings of the GG Coil leads were transmitted to Wang.

Arc Bend, Reverse Bend Dipoles (GQ, GX)

• Process Equipment Co. (PECo) is nearing completion of the GQ Cores, aiming for a July 9 inspection date with completion of the GX cores lagging by two weeks.

• We approved a change in the PECo Contract to Revision A of the drawings for both GQ and GX cores that add the shim system to the pole tip edges that makes them usable throughout the required energy range.

• At Wang NMR, seven of the eight coils for the GQ Dipoles ware now wound.


3 inch quad (QX)

• We assigned all the quadrupoles to their respective locations in the IR machine and boxed up the remaining 23 quads for the UV branch for assignment later.

Trim Quad (QT)

• Milhous Control of Virginia finally got two production units to us.

• We are now analyzing the use of the integral GC vertical corrector at several levels of quadrupole operation in anticipation of the change request required for this innovation

Sextupole (SF)

• DULY Research "core dumped" the sextupole drawings to us for checking.

Octupole (OT)

• Work started up on this magnet. We will have DULY design this magnet at a simpler level because we will probably make or assemble much of it in-house.

Beam Line and Vacuum

• The bids for the ARC Chambers were received. The technical evaluation team met and found a clear acceptable bid at a good price and acceptable delivery. A short telephone discussion cleared up some unclear points. Award now depends on internal Procurement review early next week.

• The X Chamber is coming together well in the Shop.

• The drawings for the recirculation region in front of the Cryomodule were not signed off because of some problems with the dump and some reversed antenna leads on a BPM. However the positions of stands and cartridges were checked and the girders can now be released for construction because they are verified to be identical to the 2F regions 8 girders.

• Design of the four chambers for the optical chicane region went through review of the Tera Hz port.

• Design of the regions around the wiggler continues.

WBS 11 (Optics):

We received the main mounting plates and various fabricated hardware for the optical cavities. It is in our Machine Shop, going through QC. We anticipate receiving the final QC on the OCMMS beamline parts next week. The OCMMS support stands are designed, and we are getting quotes for their fabrication and the optical tables that rest on them. The first prototype, compact 50kW beam dump is being brazed, and will be out next week. The optical cavity vacuum vessels are being procured.

We've received the prototype backplane cooled mirror holder and a test mirror. This will be checked on the mirror test stand in the near future. The designs for the mirror cassette modifications and the turning cassettes are underway.

The ultrafast laser oscillator was converted to femtosecond operation, but not without some problems. Since they were associated with the installation process, we think they are one time occurrences, and we should get dependable operation from this laser. The pulsewidth is less than 80 fs, and we are in the process of checking its other performance parameters. Some additional optics were ordered for the UV optics damage test projects.